# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

### v5

@@ 7 Building a Simulation
Parameter Ranges and Boundary Conditions
To apply TALM in realistic simulations, we constrain the variables used in the augmented persistence inequality:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t),\quad\text{where}\quad\Phi(t)=f(\Delta S_{net}^{-1},D(x,t))$
  • Energy Input $E_{\text{in}}$: $10^{-21}$ to $10^{-18}$ J per molecule per event, from UV, geothermal, or chemical sources (Schrödinger, 1944; Morowitz, 1968).
  • Energy Expenditure $E_{\text{cost}}$: $10^{-20}$ to $10^{-18}$ J to sustain reaction or structural integrity.
  • Entropy Balance $\Delta S_{net}$: Net positive when $\Delta S_{env}\sim 10^{-21}$ to $10^{-20}$ J/K exceeds local ordering $\Delta S_{local}\sim-10^{-22}$ J/K.
  • Diffusion Modifier $D(x,t)$: Confined systems (vesicles, mineral pores) exhibit $10^{-14}$ to $10^{-12}$ m2/s; bulk water approximates $10^{-9}$ m2/s.
  • Resilience Term $R_{n}$: A normalized structural redundancy factor from 0 to 1; autocatalytic or feedback-rich systems may exceed 0.5.
  • Temporal Window $t$: Simulations should explore intervals from seconds to several days (e.g., 10 to $10^{5}$ seconds), reflecting real prebiotic cycling conditions.
Boundary conditions may include:
  • Fixed or variable environmental energy sources
  • Randomized initial reactant sets or network topologies
  • Compartmental geometry or surface structure modifiers
Simulation outcomes can be interpreted in terms of the probability $P(y^{\prime}(t)\geq 0)$, and aggregated across permutations $N$ to derive $P_{\text{selective}}$ for entire planetary scenarios.
Toward a Computational Model
The TALM framework lends itself naturally to simulation. A computational model can be constructed to evaluate the persistence of reaction chains under fluctuating environmental energy input $z(t)$, internal energy transformations, and thermodynamic constraints.
@@
Future iterations may integrate entropy export ($\Delta S$), diffusion penalties, and compartmentalization effects directly into chain viability evaluations.
Baseline TALM Simulation (No Entropic or Spatial Constraints)
The following pseudocode outlines the core logic of the Thermodynamic Abiogenesis Likelihood Model (TALM) in its original, simplified form. This version models only the energy balance required to sustain a chain of prebiotic chemical reactions under a time-varying energy input $z(t)$, without accounting for entropy or diffusion penalties. Persistence is evaluated using the inequality $y(t)\geq 0$, where $y(t)$ is the system’s net residual energy after reaction execution.
1Initialize environmental energy function z(t)
2Generate a set of reaction chains
3
4For each timestep t:
5 Update environmental energy z = z(t)
6
7 For each reaction chain:
8 energy_available = z
9 Mark chain as viable
10
11 For each reaction in chain:
12 If energy_available >= required_energy:
13 Subtract required_energy from energy_available
14 Add released_energy to energy_available
15 Store any excess (optional)
16 Else:
17 Mark chain as non-viable
18 Break loop
19
20 If energy_available >= 0:
21 System persists
22 Optionally replicate or mutate
23 Else:
24 System collapses
Listing 1: Baseline TALM Simulation (No Entropy or Diffusion)
Augmented TALM Simulation with Entropy and Diffusion
This extended version of the TALM simulation incorporates entropy dynamics and diffusion-based spatial penalties into the persistence function. The augmented viability condition is given by:
@@ 24 System collapses
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t),\quad\text{where}\quad\Phi(t)=\alpha\cdot\frac{1}{\Delta S_{\text{net}}}+\beta\cdot D(x,t)$
Entropy export ($\Delta S_{\text{env}}$) and local structural order ($\Delta S_{\text{local}}$) contribute to a net entropy balance, while diffusion $D(x,t)$ reflects the dispersal potential of the system. This model better represents physical constraints in structured environments such as pores, vesicles, or hydrothermal interfaces.
1Set constants: α (entropy scale), β (diffusion scale), γ (structure factor), T (temperature)
2
3Initialize z(t)
4Generate reaction chains with:
5 - diffusion_coefficient D(x, t)
6 - resilience R_n
7
8For each timestep t:
9 Update z = z(t)
10
11 For each reaction chain:
12 energy_available = z
13 structure_index = 0
14 Mark chain as viable
15
16 For each reaction in chain:
17 If energy_available >= required_energy:
18 Subtract required_energy from energy_available
19 Add released_energy to energy_available
20 Store stored_energy (optional)
21 Increment structure_index
22 Else:
23 Mark chain as non-viable
24 Break loop
25
26 Compute entropy terms:
27 delta_S_env = exported_heat / T
28 delta_S_local = γ * structure_index
29 delta_S_net = max(delta_S_env - delta_S_local, ε)
30
31 Compute penalty:
32 phi = α / delta_S_net + β * diffusion_coefficient
33
34 Compute augmented viability:
35 y_prime = energy_available + R_n - phi
36
37 If y_prime >= 0:
38 System persists
39 Optionally replicate or mutate
40 Else:
41 System collapses
Listing 2: Augmented TALM Simulation with Entropy and Diffusion

### v7

@@ 4 Building a Simulation
4.1 Simulation Logic
The model is implemented numerically by evaluating the persistence inequality across a sequence of time-steps. Each time-step represents an interval in which the system receives environmental free energy, incurs internal energetic costs, and undergoes changes in stored energy, resilience, and entropy-related dissipation.
At each interval, the free-energy balance is computed using the unified viability function,
$y(t)=z(t)+S(t)+\sum_{i}r_{i}-\sum_{i}x_{i},$
and its thermodynamically refined form,
$y^{\prime}(t)=y(t)+R(t)-\Phi(t).$
Environmental free energy $z(t)$ is drawn from a stochastic distribution representing variable inputs such as temperature fluctuations, photochemical variability, or intermittent chemical gradients. The quantities $x_{i}$ and $r_{i}$ are set by the reaction network and represent the energetic costs and stabilising contributions specific to that configuration.
The entropic–diffusive penalty and structural resilience are evaluated directly:
$\Phi(t)=\alpha_{S}\,T\,\Delta S_{\text{local}}(t)+\beta_{D}\,\frac{RT}{D_{0}}\,D(t),\qquad R(t)=\lambda_{R}\,R(G(t)),$
ensuring consistency with the free-energy interpretation established in the theoretical framework.
Stored free energy evolves according to the discrete-time update,
$S(t+\Delta t)=S(t)+\sum_{i}s_{i}(t)-\gamma S(t),$
which is the time-stepped analogue of the continuous relation
$\frac{dS}{dt}=\sum_{i}s_{i}(t)-\gamma S(t).$
The term $\gamma S(t)$ captures generic losses such as leakage, instability, or uncaptured work.
Persistence occurs at a time-step whenever
$y^{\prime}(t)\geq 0,$
which is equivalent to a non-positive net Gibbs free-energy change over that interval. The proportion of intervals satisfying this condition across $N$ time-steps provides an estimate of the persistence frequency. These values are used to construct the free-energy landscapes presented in the results, illustrating how environmental variability, energetic costs, structural retention, and entropy production jointly determine the likelihood of sustained reaction progression.
A computational model can be constructed to evaluate the persistence of reaction chains under fluctuating environmental energy input $z(t)$, internal energy transformations, and thermodynamic constraints.
@@ 4.1 Simulation Logic
Two simulation models are proposed here in order to show the core principle behind the persistence model, and an augmented simulation designed to show how iterations of the model can incorporate entropic and spatial factors. Further iterations may integrate a number of other factors and compartmentalization effects directly into chain viability evaluations.
4.2 Parameter Ranges and Boundary Conditions
To apply the persistence model in realistic simulations, we constrain the variables used in the augmented persistence inequality:
@@ 4.2 Parameter Ranges and Boundary Conditions
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t),\quad\text{where}\quad\Phi(t)=\alpha_{S}T\,\Delta S_{\text{local}}(t)+\beta_{D}\frac{RT}{D_{0}}D(x,t)$
  • Environmental Energy Input $z(t)$: Typical fluctuations range from $10^{-21}$ to $10^{-18}\,\mathrm{J\;molecule^{-1}}$ per event, reflecting contributions from UV irradiation, geothermal cycling, or chemical redox processes [8, 20]. In the model, these values correspond directly to sampled realizations of $z(t)$.
  • Reaction Energy Requirements $x_{i}$: Energetic costs for individual reaction steps are typically $10^{-20}$ to $10^{-18}\,\mathrm{J\;molecule^{-1}}$, representing bond formation, activation barriers, or structural maintenance.
  • Reaction Energy Releases $r_{i}$: Exergonic steps return energy to the system in the same units as $x_{i}$ and $z(t)$. Typical magnitudes fall within $10^{-21}$ to $10^{-19}\,\mathrm{J\;molecule^{-1}}$, depending on reaction class.
  • Entropy Balance $\Delta S_{\mathrm{net}}$: Net entropy export is positive when environmental dissipation $\Delta S_{\mathrm{env}}\sim 10^{-21}\text{ to }10^{-20}\,\mathrm{J\,K^{-1}}$ exceeds local ordering $\Delta S_{\mathrm{local}}\sim 10^{-22}\,\mathrm{J\,K^{-1}}$. The viability penalty uses $\Delta S_{\mathrm{net}}^{-1}$ through the function $\Phi(t)$.
  • Diffusion Modifier $D(x,t)$: Effective diffusivity is environment dependent. Confined systems such as vesicles or mineral pores typically exhibit $D\sim 10^{-14}$ to $10^{-12}\,\mathrm{m^{2}\,s^{-1}}$, whereas bulk aqueous diffusion is approximately $D\sim 10^{-9}\,\mathrm{m^{2}\,s^{-1}}$. Higher diffusivity increases $\Phi(t)$.
  • Resilience Term $R_{n}$: A dimensionless parameter representing structural redundancy, buffering, or pathway multiplicity. Values range from 0 (no resilience) to 1 (high redundancy). Autocatalytic or feedback-rich networks may exhibit $R_{n}>0.5$.
  • Temporal Window $t$: Simulations commonly explore $t$ from $10$ to $10^{5}$ seconds (seconds to days), consistent with prebiotic cycling intervals such as wet-dry phases, diurnal heating, or geothermal oscillations.
Boundary conditions may include:
  • Fixed or variable environmental energy sources
  • Randomized initial reactant sets or network topologies
  • Compartmental geometry or surface structure modifiers
Simulation outcomes can be interpreted in terms of the probability $P(y^{\prime}(t)\geq 0)$, and aggregated across permutations $N$ to derive $P_{\text{selective}}$ for differing alternative scenarios.
4.2.1 Simulation defaults and reproducibility.
Unless otherwise noted, simulations employ $\bar{z}\in[10^{-21},10^{-19}]$ J molecule-1, $\sigma_{z}=0.2\,\bar{z}$, $D\in\{10^{-14},10^{-12},10^{-9}\}$ m2s-1, $\gamma\in\{10^{-6},10^{-4}\}$ s-1, $\alpha_{S}=\beta_{D}=1$, and $R_{n}=RT\,\mathcal{R}(G)$ at $T=330$ K. Each parameter pair is evaluated using $N_{\mathrm{trials}}=10^{3}$ across $N_{\mathrm{cycles}}=10^{3}$ iterations. The resulting phase boundaries remain stable under $\pm 20\%$ variation, indicating robust qualitative behavior.
4.3 Baseline Simulation (No Entropic or Spatial Constraints)
The persistence inequality was implemented as a stochastic simulation to evaluate the fraction of viable reaction chains under varying environmental inputs. Each simulation represents an open system experiencing discrete energy intervals, during which the environmental input $z(t)$ fluctuates as a random variable drawn from a normal distribution with mean $\bar{z}$ and standard deviation $\sigma_{z}$. Stored energy $S(t)$ accumulates and dissipates across intervals according to $\frac{dS}{dt}=\sum s_{i}-\gamma S(t)$. The algorithmic form is given below.
Algorithm 1 Baseline Persistence Simulation (No Entropy or Diffusion)
[…]
3:for each time-step $t$ do
4:  $z\leftarrow$ SampleEnvironmentalEnergy($t$)
5:  for each reaction chain $\mathcal{C}$ do
6:   $E\leftarrow z+S$ $\triangleright$ Initial available free energy
7:   $\mathcal{C}.\text{viable}\leftarrow\text{True}$
8:   for each reaction $i$ in $\mathcal{C}$ do
9:     $(x_{i},r_{i},s_{i})\leftarrow$ ReactionParameters(ii)
10:     if $E<x_{i}$ then
11:      $\mathcal{C}.\text{viable}\leftarrow\text{False}$
12:      break
13:     end if
14:     $E\leftarrow E-x_{i}+r_{i}$
15:     $S\leftarrow S+s_{i}$
16:   end for
17:   $\mathcal{P}\leftarrow\mathcal{P}\cup\{(t,\mathcal{C},\mathcal{C}.\text{viable})\}$
18:  end for
19:  $S\leftarrow S-\gamma S$ $\triangleright$ Loss of stored free energy
20:end for
21:return $\mathcal{P}$
Figure 2: Phase boundary between persistence and collapse predicted by the framework. The heatmap shows the fraction of simulated trials in which $y^{\prime}(t)\geq 0$ across multiple environmental intervals, plotted against the mean environmental energy input $\bar{z}$ and its variability $\sigma_{z}$. Higher $\bar{z}$ and lower variability favour persistence, demonstrating a clear transition between viable and non-viable regimes.
The baseline simulation captures persistence dynamics driven solely by variable energy input and storage.
4.4 Augmented Simulation (Entropic and Spatial Constraints)
To examine how spatial and entropic factors influence viability, an augmented implementation introduces the extended penalty term $\Phi(t)$ and a resilience parameter $R_{n}$ that modulates the effect of structural redundancy.
The augmentation introduces an entropic-diffusive penalty $\Phi(t)=\alpha_{S}T\,\Delta S_{\text{local}}(t)+\beta_{D}\frac{RT}{D_{0}}D(x,t)$ and a resilience factor $R_{n}$. This preserves the baseline simulation structure while allowing persistence boundaries to shift with entropy gradients and diffusion rates.
Algorithm 2 Augmented Persistence Simulation with Entropic and Diffusive Modifiers
[…]
3:for each time-step $t$ do
4:  $z\leftarrow$ SampleEnvironmentalEnergy($t$)
5:  for each reaction chain $\mathcal{C}$ do
6:   $E\leftarrow z+S$
7:   $\mathcal{C}.\text{viable}\leftarrow\text{True}$
8:   for each reaction $i$ in $\mathcal{C}$ do
9:     $(x_{i},r_{i},s_{i})\leftarrow$ ReactionParameters(ii)
10:     if $E<x_{i}$ then
11:      $\mathcal{C}.\text{viable}\leftarrow\text{False}$
12:      break
13:     end if
14:     $E\leftarrow E-x_{i}+r_{i}$
15:     $S\leftarrow S+s_{i}$
16:   end for
17:   if $\mathcal{C}.\text{viable}=\text{False}$ then
18:     $\mathcal{P}\leftarrow\mathcal{P}\cup\{(t,\mathcal{C},\text{False})\}$
19:     continue
20:   end if
21:   $\Delta S_{\text{loc}}\leftarrow$ ComputeLocalEntropyChange($\mathcal{C},t$)
22:   $D\leftarrow$ ComputeDiffusivity($\mathcal{C},t$)
23:   $R\leftarrow$ ComputeResilience($\mathcal{C}$)
24:   $\Phi\leftarrow\alpha_{S}\cdot T\cdot\Delta S_{\text{loc}}\;+\;\beta_{D}\cdot(R_{g}T/D_{0})\cdot D$
25:   $y_{\text{base}}\leftarrow E$
26:   $y_{\text{aug}}\leftarrow y_{\text{base}}+R-\Phi$
27:   if $y_{\text{aug}}\geq 0$ then
28:     $\mathcal{C}.\text{viable}\leftarrow\text{True}$
29:   else
30:     $\mathcal{C}.\text{viable}\leftarrow\text{False}$
31:   end if
32:   $\mathcal{P}\leftarrow\mathcal{P}\cup\{(t,\mathcal{C},\mathcal{C}.\text{viable})\}$
33:  end for
34:  $S\leftarrow S-\gamma S$
35:end for
36:return $\mathcal{P}$
Figure 3: Augmented persistence boundary including the entropic-diffusive penalty $\Phi(t)=\alpha/\Delta S_{\text{net}}+\beta D$. The heatmap shows the fraction of trials with $y^{\prime}(t)\geq 0$ across intervals as a function of mean environmental input $\bar{z}$ and diffusion coefficient $D$ at fixed variability $\sigma_{z}$. Increasing $D$ raises the penalty and shifts the boundary to higher $\bar{z}$, illustrating how spatial dispersion can suppress persistence even under the same mean input.
4.4.1 Persistence Landscape Construction
To evaluate how variations in environmental energy availability influence the likelihood of reaction persistence, the model computes a persistence landscape across a range of parameter combinations. Each point in this landscape represents the proportion of time-steps for which the augmented persistence condition $y^{\prime}(t)\geq 0$ is satisfied under a particular environmental configuration. The procedure samples environmental parameters, runs multiple independent simulations using either the baseline or augmented update rules, and records the frequency of viable intervals. By repeating this process across a grid of parameter values, the resulting matrix forms a quantitative representation of the persistence landscape, which serves as the basis for the heatmaps and phase regions shown in the results. Algorithm 3 summarises the computation of this landscape.
Algorithm 3 Persistence Landscape Generation
1:Set of environmental parameter combinations $\{\theta_{k}\}$
2:Number of time-steps $N_{\text{steps}}$
3:Number of independent runs per parameter set $N_{\text{runs}}$
4:Simulation mode (Baseline or Augmented)
5:Initialise landscape matrix $L\leftarrow\emptyset$
6:for each parameter set $\theta_{k}$ in $\{\theta_{k}\}$ do
7:  $\text{persistent\_intervals}\leftarrow 0$
8:  $\text{total\_intervals}\leftarrow 0$
9:  for $r\leftarrow 1$ to $N_{\text{runs}}$ do
10:   ConfigureEnvironment($\theta_{k}$)
11:   Initialise stored free energy $S\leftarrow 0$
12:   for $t\leftarrow 1$ to $N_{\text{steps}}$ do
13:     if SimulationMode = Baseline then
14:      $\mathcal{P}_{t}\leftarrow$ RunBaselineStep($t$, $S$)
15:     else
16:      $\mathcal{P}_{t}\leftarrow$ RunAugmentedStep($t$, $S$)
17:     end if
18:     for each record $(t,\mathcal{C},\mathrm{viable})$ in $\mathcal{P}_{t}$ do
19:      $\text{total\_intervals}\leftarrow\text{total\_intervals}+1$
20:      if $\mathrm{viable}=\text{True}$ then
21:        $\text{persistent\_intervals}\leftarrow\text{persistent\_intervals}+1$
22:      end if
23:     end for
24:   end for
25:  end for
26:  $f_{\text{persist}}(\theta_{k})\leftarrow\text{persistent\_intervals}/\text{total\_intervals}$
27:  Store $f_{\text{persist}}(\theta_{k})$ in $L$ at the coordinates corresponding to $\theta_{k}$
28:end for
29:return landscape matrix $L$
Summary
Taken together, the three simulation procedures provide a coherent computational counterpart to the theoretical persistence framework. The baseline algorithm evaluates viability under fluctuating environmental free-energy inputs, while the augmented algorithm incorporates the additional entropic, diffusive, and structural contributions that shape the free-energy balance over time. Algorithm 3 extends these procedures to construct persistence landscapes, enabling systematic exploration of how different environmental and molecular parameters influence the frequency of intervals in which the condition $y^{\prime}(t)\geq 0$ is satisfied. These simulations therefore translate the analytical model into quantitative predictions, allowing the behaviour of diverse hypothetical reaction networks to be compared across parameter ranges and providing an empirical foundation for the persistence maps presented in the results. In subsequent sections, this computational foundation is generalized into the Thermodynamic Abiogenesis Likelihood Model (TALM), which extends these persistence criteria to broader chemical and planetary contexts.
5 Thermodynamic Abiogenesis Likelihood Model (TALM)
The Thermodynamic Abiogenesis Likelihood Model (TALM) applies the persistence inequality and its modulators as an analytical framework for evaluating viability across chemical ensembles and planetary environments. It generalizes the persistence inequality into a probabilistic framework that incorporates entropic transformations, diffusion constraints, and spatial organization as modulators of molecular persistence.
5.1 Entropy-Driven Exploration of Chemical Space
TALM adopts the view that in open, far-from-equilibrium systems, entropy production enables exploration of configuration space: variable environmental energy fluxes increase global entropy while permitting localized entropy reductions associated with transient structures and reaction networks. Stochastic configurations generated in this way are then filtered by the reaction viability inequality, linking entropy-driven exploration to thermodynamic selection.
5.2 Reaction Viability in Structured Environments
TALM integrates the augmented viability function:
$y^{\prime}(t)=y(t)+R(t)-\Phi(t)$ (5)
Here the components are interpreted as follows:
  • $y(t)$ is the underlying free-energy balance incorporating environmental input, stored free energy, and reaction-associated energetic terms
  • $R(t)$ is the resilience contribution associated with structural or contextual stability
  • $\Phi(t)$ is the entropic–diffusive penalty associated with maintaining local organization
The penalty term and resilience contribution are scaled using two constants:
  • $\alpha_{S}$ scales the entropy-related cost of maintaining local order
  • $\beta_{D}$ scales the diffusive cost associated with spatial dispersion
The entropic–diffusive penalty is therefore written as
$\Phi(t)=\alpha_{S}\,T\,\Delta S_{\text{local}}(t)+\beta_{D}\,\frac{RT}{D_{0}}\,D(x,t),$
which captures the free-energy cost of counteracting thermal disorder and diffusion in non-equilibrium environments. This formulation provides a dimensionally consistent representation aligned with the augmented persistence model.
5.3 Likelihood Formulation
The likelihood that at least one interval supports persistence is
$P_{\text{selective}}=1-\big(1-P(y^{\prime}(t)\geq 0)\big)^{N_{\text{eff}}},$ (6)
where $P(y^{\prime}(t)\geq 0)$ is the probability that the free-energy balance is favourable during a sampled interval, and $N_{\text{eff}}$ is the effective number of independent instances. In spatially or temporally correlated settings, independence is reduced such that
$N_{\text{eff}}=\frac{N}{\,1+2\sum_{\tau\geq 1}\rho(\tau)\,}$,
with $\rho(\tau)$ representing the autocorrelation of viability across successive intervals. This correction accounts for partial dependence between samples in confined or resource-limited environments.
5.4 Comparative Context: Earth and Icy Moons
To illustrate the applied potential of the TALM framework, we consider its implications for three planetary environments: early Earth, Europa, and Enceladus. These contexts exhibit the chemical diversity, compartmentalisation opportunities, and variable energy inputs relevant to persistence dynamics. These examples are illustrative rather than quantitative and demonstrate how the persistence inequality may be used to compare environmental favourability rather than to provide specific predictions.
Early Earth featured diurnal heating, geothermal activity, and intermittent hydration intervals, producing confined microenvironments (e.g., clay layers, mineral pores) with reduced diffusivity and variable free-energy supply, factors that increase the probability of $y^{\prime}(t)\geq 0$ within TALM.
Europa may possess hydrothermal activity beneath an icy shell, with tidal flexing driving sustained energy input. Although lacking surface cycling, the combination of geothermal gradients and mineral interfaces could support moderate $P_{\text{selective}}$ values, particularly in porous or compartmentalised regions.
Enceladus shows evidence of internal heating and organic-rich plumes, suggesting transient microenvironments with spatial structure and chemical gradients. TALM predicts that even short-lived favorable intervals could support molecular persistence where thermal or chemical pulses align with free-energy requirements.
These examples illustrate how TALM provides a qualitative means of comparing abiogenesis potential across planetary settings.
5.5 Urability as a Prebiotic Filter
Just as habitability describes environments where life can be sustained, urability refers to environments capable of supporting the origin of life. It characterizes the prebiotic window in which physical and chemical conditions—such as compartmentalisation, accessible substrates, and non-equilibrium free-energy flux—enable the emergence of molecular complexity prior to heredity or replication [21].
Within the TALM framework, urability functions as a pre-filter for identifying planetary settings where preconditions for abiogenesis are plausible. TALM then evaluates whether reaction networks in those settings satisfy the persistence criterion with sufficient frequency. Together, urability and TALM provide a complementary two-tier approach to assessing the likelihood of abiogenesis.
6 Illustration of the Model
6.1 Reaction chain model
6.2 Energy dynamics over time
Figure 4 illustrates energy dynamics over time, showing the interaction between stored energy and energy used in reactions.
Figure 4: The energy–time plot illustrates how environmental energy input and internal energy storage jointly determine the viability of a reaction sequence across fluctuating conditions.
6.3 Numerical Example: Reaction Chain Viability
6.3.1 Reaction Chain Parameters
We define the following reactions:
Reaction $x_{i}$ (Required) $r_{i}$ (Released) $s_{i}$ (Stored)
$R_{1}$ 8 2 1
$R_{2}$ 5 3 2
$R_{3}$ 7 1 1
6.3.2 Case 1: Sufficient Environmental Energy ($z(t)=10$
  • $R_{1}$: $z=10\rightarrow z-x_{1}+r_{1}=10-8+2=4\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=4\rightarrow z-x_{2}+r_{2}=4-5+3=2\rightarrow$ Store $s_{2}=2$
Result: System collapses at step 3.
6.3.3 Case 2: Higher Energy Input ($z(t)=15$)
  • $R_{1}$: $z=15\rightarrow 15-8+2=9\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=9\rightarrow 9-5+3=7\rightarrow$ Store $s_{2}=2$
  • $R_{3}$: $z=7\rightarrow 7-7+1=1\rightarrow$ Store $s_{3}=1$
6.4 Illustrative Example: Thermal Cycling on Early Earth
To illustrate the applicability of the persistence framework under real environmental conditions, consider empirical measurements by Deamer and Damer (2020), who demonstrated that dehydration-rehydration intervals in hydrothermal fields drive non-enzymatic polymerization of nucleotide-like molecules under prebiotic conditions by recorded daily temperature fluctuations ranging from 30-90°C in surface hydrothermal pools at Kamchatka[22] [10]. These fluctuations correspond to molecular-scale energy variations on the order of $10^{-20}$ to $10^{-18}$ joules per molecule making ideal testbeds for modeling $z(t)$ the time-dependent environmental energy input.
$z(t)=z_{\max}\cdot\sin(\omega t)+z_{\text{offset}}$
During daytime, energy peaks may support energetically demanding reactions ($\sum x_{i}$ is high), while nighttime introduces an energy deficit. Only systems with sufficient internal storage $S(t)$, or inherently low energy requirements $x_{i}$, will remain viable during low-input phases.
For example, let a reaction system require $x_{i}=5$ energy units to maintain structural viability and release $r_{i}=1$. If $z(t)$ peaks at 6 and falls to 1 at night, then persistence through a full cycle is only possible for systems that can retain at least $S(t)\geq 3$ units of stored energy from the high-input phase.
Figure 5: Simulated thermal cycling and entropy export in early Earth conditions. The solid curve represents energy input $z(t)$ over a 48-hour period, modeled as a sinusoidal day-night cycle with baseline energy. The dashed curve shows entropy exported to the environment, $\Delta S_{env}(t)$, phase-shifted to reflect thermal dissipation lag. The shaded region indicates the viable persistence zone, where energy input exceeds the threshold required for reaction continuity. This illustrates how persistent systems must not only align with fluctuating $z(t)$, but also export sufficient entropy to remain thermodynamically viable.
This scenario demonstrates the central principle behind the persistence model: differential energy storage enables selective retention. Systems capable of buffering or storing energy during abundance become preferentially stable across environmental fluctuations. Thermal cycling, when combined with compartmentalization and diffusion-limiting substrates, thus constitutes a plausible natural engine for prebiotic selection.
Furthermore, such persistence would only be thermodynamically viable if local decreases in entropy, associated with maintaining molecular structure across intervals, are offset by corresponding entropy export to the environment. In this way, day-night thermal cycling could support transient self-organization that satisfies the second law, especially in compartmentalized or surface-adhered systems where diffusion is constrained. This type of modeling could be extended to other environments with time-variable or localized energy gradients, such as hydrothermal vents.
6.5 Illustrative Application to Planetary Environments
Although this work focuses primarily on early Earth and hydrothermal field settings, the framework is not restricted to terrestrial conditions. In principle, any planetary or lunar environment that can be characterized in terms of (i) energy input statistics $z(t)$, (ii) chemical diversity and availability of reactive substrates, and (iii) the presence or absence of diffusion-limiting structures can be mapped onto the persistence inequality and its augmented form.
A practical application would proceed in two stages. First, an environment would be assessed for urability in the sense of Deamer and colleagues [21]: whether it plausibly offers fluctuating energy sources, suitable solvents, and some means of spatial confinement. Second, for environments that meet these baseline criteria, the persistence model could be utilized to sample ensembles of reaction networks under estimated $z(t)$ profiles and diffusion constraints, yielding comparative values of $P_{\text{selective}}$ across candidate worlds. This would not provide a literal probability that life exists, but rather a physically grounded ranking of environments by their propensity to support persistent, energy-compatible reaction networks.
A full implementation of such a planetary survey lies beyond the scope of the present paper, but the framework outlined here is expressly designed to permit this type of extension once appropriate environmental and geochemical constraints are available.